# ThinkTank: A Benchmark for Cognitive Intelligence in Pre-trained Language Models

### v1

@@ Abstract
Recent studies show evidence for emergent cognitive abilities in Large Pre-trained Language Models (PLMs). The increasing cognitive alignment of these models has made them candidates for cognitive science theories. Prior research into the emergent cognitive abilities of PLMs has been path independent to model training, i.e., only looks at the final model weights and not the intermediate steps. However, building plausible models of human cognition using PLMs requires also aligning their performance during training to the developmental trajectories of children’s thinking. To catalyze this effort, we build ThinkTank: a comprehensive benchmark of tasks for PLMs, guided by psychometric tests of human intelligence. We investigate the cognitive alignment of popular PLMs and evaluate each of their available intermediate and final training steps on a suite of cognitively relevant tasks: Numerical ability, Linguistic abilities, Conceptual understanding, and Fluid reasoning. We find a striking regularity: regardless of model size, the developmental trajectories of PLMs always show a window of maximal alignment to human cognitive development. Before that window, training appears to endow “blank slate” models with the requisite structure to be poised to rapidly learn from experience. After that window, training appears to serve the engineering goal of reducing loss but not the scientific goal of increasing alignment with human cognition.
@@ 1 Introduction
Figure 1: A list of cognitive intelligence tasks in ThinkTank. (A): Numeric intelligence, (B): Linguistic intelligence, (C): Conceptual intelligence, (D) Fluid intelligence.
@@
Large Pre-trained Language Models (PLMs) like Google’s Gemini geminiteam2023gemini, Meta’s LLaMA 2 touvron2023llama, and OpenAI’s GPT 4 openai2023gpt4 show human-level or even super-human performance on many cognitive performance tasks. This is true in domains such as mathematical reasoning shah2023numeric; ahn2024large, language comprehension warstadt2020blimp; ye2023comprehensive; koubaa2023gpt, concept understanding cog_sci_typicallity, and analogical reasoning webb2023emergent; hu2023context, contributing to the hype of claims of reaching Artificial General Intelligence (AGI).
@@
Such claims deserve to be scrutinized. Human intelligence is multi-faceted. Psychometric measures of intelligence include multiple subtests spanning a range of abilities, including mathematical thinking, language comprehension, spatial thinking, fluid reasoning, and conceptual understanding snow1984topography; carroll1993human; sternberg2000handbook; mcgrew2009chc; haier2023neuroscience. One contribution of our work is ThinkTank: a comprehensive suite of tasks to evaluate the psychometrically defined “intelligence” of PLMs. These tasks are representative assessments of the different facets of human intelligence, modified for the textual modality required for PLMs. The only domain that is noticeably missing is spatial thinking, which is harder to operationalize in text-based models.
@@
A second goal of our work is to move beyond evaluating the cognitive alignment of PLMs and to also consider their developmental alignment. The claim that the final model state of a PLM approximates adult performance leaves open the question of the path by which it arrived there. Ideally, the model’s performance improvements over training also track the progression of cognitive abilities over development elman1996rethinking; bengio2009curriculum. This potential parallelism would be stronger evidence for PLMs as cognitive science models. Researchers are increasingly asking this question by building PLMs trained on a developmentally plausible corpus of child-directed speech, transcribed dialogue, and children’s literature bhardwaj2024pretraining; huebner-etal-2021-babyberta; conll-2023-babylm.
@@
We ask the question of developmental alignment in a theoretically important way: Is the cognitive alignment of PLMs achieved in a path-independent or path-dependent manner? Prior studies focusing on the cognitive alignment of PLMs have only established path independence: that models at the end of training approximate adult performance across a range of domains. Here, we also evaluate path dependence: Do the performance improvements of PLMs over training track the growth of these abilities in children over development holyoak1984development? We ask this question for models of different sizes and track their developmental alignment over millions and billions of training tokens. If path independence holds, this opens up applications that we detail in the conclusion.
@@
  • ThinkTank: A first-of-its-kind benchmark of human-like psychometric intelligence tests in PLMs. This benchmark evaluates numeric, linguistic, conceptual, and fluid intelligence. We propose this benchmark to be a prerequisite to using PLMs for cognitive modeling.
  • Developmental trajectories in LLM pre-training and scaling: Previous studies have only evaluated the final training checkpoints of PLMs for their cognitive plausibility, and have neglected the question of developmental trajectories. We consider a variety of PLM architectures and evaluate each of their available intermediate and final training steps on ThinkTank.
Table 1: Summary of assessments in ThinkTank.
@@
Numeric Abilities Magnitude Comparison Effects shah2023numeric (cc by 4.0)
Linguistic Abilities BLiMP warstadt2020blimp (cc by 4.0)
Concept Understanding Typicality Effects cog_sci_typicallity; castro21 (cc by 4.0)
Fluid reasoning Raven’s Progressive Matrices hu2023context (cc by 4.0)
@@ 2.1 Psychometric theories of intelligence
Rather than choose cognitive abilities in an ad hoc fashion, we look to psychometric theories of intelligence while building ThinkTank sternberg2000handbook. These theories begin with a large number of subtests of a range of cognitive abilities that are given to a large sample of individuals. The matrix of correlations between pairs of subtests is reduced using factor-analytic techniques to identify a smaller number of latent factors which are “interpreted” as a theory of the core domains of intelligence kaufman2019psychometric. Despite popular attention to “general intelligence” and the latent factor g, there is a long history of theories positing that intelligence is composed of multiple domain-specific abilities. An important, early domain-specific theory of intelligence, thurstone1938primary, included seven “primary abilities”. The most widespread psychometric theory of intelligence today, the Cattell-Horn-Carrol (CHC) theory carroll1993human; mcgrew2009chc, includes among its “broad” abilities quantitative knowledge, reading and writing ability, fluid reasoning, and “comprehension” knowledge (a subcomponent of which is conceptual understanding). We evaluate the cognitive and developmental alignment of PLMs along these four abilities.
@@ 2.2 Emergent cognitive abilities in Language Models
Recently, the performance of language models has improved dramatically as they have increased in size from millions to billions of parameters, trained on larger corpora, and further tuned in novel ways (instruction tuned, RLHF). This has led to more researchers advocating for the use of PLMs as cognitive models piantadosi2023modern; warstadt2024artificial. Increasing the number of parameters of the models has proven evidence for Emergent Abilities wei2022emergent, which are abilities that cannot be predicted by extrapolating from the performance of smaller models. Emergent abilities have been observed in a variety of task types such as multi-task language understanding hendrycks2021measuring, grounded conceptual mapping patel2022mapping, truthfulness lin2021truthfulqa, etc. In recent works, hoffmann2022training; biderman2023pythia show the benefits of training a model for more tokens on problem-solving wei2022chain, common-sense reasoning sakaguchi2021winogrande, arithmetic abilities biderman2023pythia, and linguistic phenomena paperno2016lambada. Although the presence of emergent abilities extends to cognitive science domains wei2022chain; goertzel2023generative; hagendorff2023machine, prior studies have been piecemeal in their approach and have failed to (1) consider multiple cognitive abilities as specified by theories of psychometric intelligence and (2) move beyond cognitive alignment to also evaluate the developmental alignment of PLMs over training. Thus, we evaluate the cognitive and developmental alignment of PLMs from two perspectives, that of psychometric theories of intelligence and cognitive science.
3 ThinkTank: A suite of psychometric intelligence tasks
We introduce ThinkTank, a set of assessments that benchmark PLMs across four abilities of psychometric intelligence. The structure of ThinkTank is summarized in Table 1 along with the licensing details for public use. All ThinkTank tasks and corresponding datasets have low ethical risks and none expose sensitive information. Additionally, we obtain approval for the authors of each dataset for the ThinkTank benchmark. The details of each assessment and their respective operationalization are given below. ThinkTank will be added to a publically available unified language model testing framework, titled lm-evaluation-harness eval-harness, to support the evaluation of future models of psychometric intelligence assessments.
@@ 3.1 Numeric abilities
The question of how humans understand symbolic numbers has been investigated by cognitive scientists for more than half a century. These studies show that people map number symbols to a mental number line (MNL, Figure 2) with a log-compressed psychophysical scale moyerTimeRequiredJudgements1967.
Prior research on the numerical abilities of PLMs has focused on improving performance on application-driven tasks that require numerical skills in the context of arithmetic equations and word problems awp_1; awp_2; yuan2023large, exact facts  exact_facts, and measurement estimation  approx_measurement. However, these tasks fail to directly implicate the key cognitive construct underlying human numerical understanding, the recruitment of a compressed MNL.
In a recent study, shah2023numeric found evidence for a human-like MNL in various PLMs. They show that despite not having explicit neural circuitry to represent numbers, through experience (i.e., vast amounts of training data), PLMs show human-like performance profiles and learn human-like representations for numerical concepts.
We follow shah2023numeric and look for the two behavioral signatures of a compressed number line representation, the distance effect and the ratio effect. In humans, these are defined as:
  • Distance effect (refer to Figure 1A - top): The greater the distance $|x-y|$ between two numbers $x$ and $y$, the faster they are compared, i.e., the greater (or lesser) number is identified MOYER1967.
  • Ratio effect (refer to Figure 1A - bottom): The time to compare two numbers $x$ and $y$ is a decreasing function of the ratio of the larger number over the smaller number $\frac{max(x,y)}{min(x,y)}$ halberdaIndividualDifferencesNonverbal2008.
These effects can be mapped to language models by adopting the following linking hypothesis: the greater the cosine similarity of two number representations in a PLM, the longer it takes to discriminate them, i.e., to which one is greater (or lesser).
Operationalization: We used the same protocol as shah2023numeric. For each effect, we test the three formats of number representations of PLMs (mixed-case number words, lower-case number words, and digits). We present the $R^{2}$ values for the Distance and Ratio effects, which are averaged across each input representation. The $R^{2}$ values for the distance effect in PLMs are obtained by fitting a linear function predicting the cosine similarity of $x$ and $y$ from their distance $|x-y|$. Similarly, $R^{2}$ values for the ratio effect in PLMs are obtained by fitting a negative exponential function predicting the normalized cosine similarity of $x$ and $y$ from their ratio $\frac{\max(x,y)}{\min(x,y)}$. The results for our replication of the other experiments conducted by shah2023numeric are in Appendix B.1. Note: This task requires access to the latent representations of models.
@@ 3.2 Linguistic abilities
Language (or verbal) ability is a central component of human cognition and cognitive neuroscience hagoort2019human. At the dawn of the cognitive revolution, it was conceptualized as a largely innate ability, and language acquisition was understood as requiring relative learning from experience chomsky2014aspects; fodor1985precis. More recently, cognitive developmentalists have shown that infants can learn language through exposure to the statistical regularities of the linguistic environment saffran1996statistical; siegelman2020statistical. These findings have been modeled using multi-layer perceptrons elman1996rethinking and, more recently, PLMs lake2023word.
We use BLiMP: The Benchmark of Linguistic Minimal Pairs for English warstadt2020blimp to evaluate the linguistic abilities of each PLM under consideration. BLiMP consists of 67 datasets of 1000 pairs of minimally different sentences which vary in grammatical acceptability and span 12 phenomena at three levels of language: morphology, syntax, and semantics. The 12 phenomena are described in Appendix B.2. Each pair consists of one acceptable sentence and one unacceptable sentence which otherwise differ minimally. BLiMP evaluates the models by measuring if they assign a higher probability to the acceptable vs. unacceptable sentence of each pair. Figure 1B shows two examples of minimal pairs.
Operationalization: We use the LM-eval-harness eval-harness benchmarking suite to test our models on the BLiMP tasks. We evaluate if the model assigns a higher sequential probability to the acceptable sentence. Note: This requires models that can generate probabilities of tokens.
@@ 3.3 Concept understanding
On encountering a new stimulus, humans categorize it – assign it to a known concept – in order to make inferences about its unobservable properties murphy2002big. Not all members of a category are equal rosch75. Rather, some members (e.g., pigeons) are more typical of a category (e.g., Birds) than other members (e.g., ostrich). This phenomenon, known as the Typicality Effect, is a central feature of human categorization lakoff2008women.
Typicality gradients in humans can be measured using the production task, where participants are given a category label (e.g., bird) and asked to list as many members of the category as they can in a limited time battig1969category; castro21; van2004category. The typicality of an item is defined as the proportion of participants who produce it.
Language models have shown some evidence of human-like typicality gradients. heyman2019prediction used word2vec embeddings to predict the category typicality norms of de2008exemplar. More recent work by misra2021language and bhatia2022transformer has looked at correlations of PLMs like BERT, RoBERTa, and GPT-2 to the rosch75 typicality norms for ten categories. cog_sci_typicallity performed the most comprehensive study of the alignment of concept understanding in the latent representations of PLMs. We expand upon their task setup to evaluate human-like concept understanding in the PLMs that are the focus here.
@@
  • Surprisal values: For each member in a category, the probability of the sequence a "member" (eg. pigeon) is a "category" (eg. bird). This method requires access to the probability of each token in a sequence.
  • Prompting: Prompt the models with the following design: Guidelines, Query, and Options. The Guideline highlights the task of re-ranking the members given in the Options based on appropriateness with the Query. The Query consists of the in-filling task: A   is a [category name]. The Options are each of the possible members of the category. Note: Given the complexity of the prompting guidelines, usable outputs are only obtained from models that are larger than 30 billion parameters.
@@
Note: Under the closeness judgment protocol, our experiments fail to match up to the performance of the models used by cog_sci_typicallity. This is because our choice of open-source models only provides token representations, on which we later perform an aggregation operation. This aggregation operation leads to a loss of information. In contrast, cog_sci_typicallity use sentence-transformer models sentence-transformers, which provide singular latent representation for longer text. This variation in experimentation leads to the difference in alignment scores.

### v3

# Development of Cognitive Intelligence in Pre-trained Language Models

@@ Abstract
Recent studies show evidence for emergent cognitive abilities in Large Pre-trained Language Models (PLMs). The increasing cognitive alignment of these models has made them candidates for cognitive science theories. Prior research into the emergent cognitive abilities of PLMs has largely been path independent to model training, i.e., has focused on the final model weights and not the intermediate steps. However, building plausible models of human cognition using PLMs would benefit from considering the developmental alignment of their performance during training to the trajectories of children’s thinking. Guided by psychometric tests of human intelligence, we choose four sets of tasks to investigate the alignment of ten popular families of PLMs and evaluate their available intermediate and final training steps. These tasks are Numerical ability, Linguistic abilities, Conceptual understanding, and Fluid reasoning. We find a striking regularity: regardless of model size, the developmental trajectories of PLMs consistently exhibit a window of maximal alignment to human cognitive development. Before that window, training appears to endow “blank slate” models with the requisite structure to be poised to rapidly learn from experience. After that window, training appears to serve the engineering goal of reducing loss but not the scientific goal of increasing alignment with human cognition.
@@ 1 Introduction
Figure 1: A list of cognitive intelligence tasks under consideration.
@@
Large Pre-trained Language Models (PLMs) like Google’s Gemini Team et al. (2023), Meta’s LLaMA 2 Touvron et al. (2023), and OpenAI’s GPT 4 OpenAI (2023a) show human-level or even super-human performance on many cognitive performance tasks. This is true in domains such as mathematical reasoning Shah et al. (2023); Ahn et al. (2024), language comprehension Warstadt et al. (2020); Ye et al. (2023); Koubaa (2023), concept understanding Vemuri et al. (2024), and analogical reasoning Webb et al. (2023); Hu et al. (2023). These successes have contributed to the hype of reaching Artificial General Intelligence (AGI).
Such claims deserve to be scrutinized. There is a massive disparity between the training data scale of PLMs and humans. However, PLMs unintentionally acquire human performance characteristics from the corpora they are trained on, through residues of the values, beliefs, and biases of the authors of the texts Pellert et al. (2024).
@@
We approach the human alignment of PLMs by grounding their evaluation in frameworks for psychometric intelligence. Psychometric measures of intelligence include multiple subtests spanning a range of abilities, including mathematical thinking, language comprehension, spatial thinking, and fluid reasoning Snow et al. (1984); Carroll (1993); Sternberg (2000); McGrew (2009); Haier (2023). In this work, we choose representative assessments of different facets of human intelligence, modified for the required textual modality, to evaluate the cognitive alignment of PLMs.
@@
A second goal of our work is to move beyond cognitive alignment to also evaluate the developmental alignment of PLMs. The claim that the final model state of a PLM approximates adult performance leaves open the question of the path by which it arrived there. Ideally, the model’s performance improvements over training should also track the progression of cognitive abilities over development Elman (1996); Bengio et al. (2009). This potential parallelism would be stronger evidence for PLMs as cognitive science models. Researchers are increasingly addressing this question by building PLMs trained on a developmentally plausible corpus of child-directed speech, transcribed dialogue, and children’s literature Huebner et al. (2021); Warstadt et al. (2023); Bhardwaj et al. (2024).
@@
We ask the question of developmental alignment in a theoretically important way: Is the cognitive alignment of PLMs achieved in a path-independent or path-dependent manner? Prior studies focusing on the cognitive alignment of PLMs have only established path independence: that models at the end of training approximate adult performance across a range of domains. Here, we also evaluate path dependence: Do the performance improvements of PLMs over training track the growth of these abilities in children over development? We ask this question for models of different sizes and track their developmental alignment over millions and billions of training tokens.
@@
  • Cognitive Modelling using AI: We test the appropriateness of PLMs for cognitive modeling by evaluating whether their performance profiles match those of adults.
  • Developmental trajectories in LLM pre-training and scaling: Previous studies have largely focused on evaluating the final training checkpoints of PLMs for their cognitive plausibility, and have neglected the question of developmental trajectories. Here, we also ask: Can PLMs be used to model developmental trajectories of children’s thinking despite the training data scale mismatch?
  • Representative tasks: We choose representative tasks of human cognition taken from psychometric tests of intelligence tests. These tasks evaluate numeric, linguistic, conceptual, and fluid intelligence. We propose these to be a prerequisite for considering PLMs as cognitive science models.
@@
Numeric Abilities Magnitude Comparison Effects Shah et al. (2023) (cc by 4.0)
Linguistic Abilities BLiMP Warstadt et al. (2020) (cc by 4.0)
Concept Understanding Typicality Effects Vemuri et al. (2024); Castro et al. (2021) (cc by 4.0)
Fluid reasoning Raven’s Progressive Matrices Hu et al. (2023) (cc by 4.0)
Table 1: Summary of assessments.
@@ 2.1 Psychometric theories of intelligence
Previous studies of the human alignment of ML models have typically looked at singular dimensions, such as numeric abilities Zhuang et al. (2023); Fang et al. (2024). Rather than choose cognitive abilities in a piecemeal fashion, we look to psychometric theories of intelligence for guidance Sternberg (2000). These theories distill performance on a large number of subtests into a small number of latent factors. Despite popular attention to “general intelligence” and the latent factor g, there is a long history of theories positing that intelligence is composed of multiple domain-specific abilities. An important, early domain-specific theory of intelligence included seven “primary abilities” Thurstone (1938). The most widespread psychometric theory of intelligence today, the Cattell-Horn-Carrol (CHC) theory Carroll (1993); McGrew (2009), includes among its “broad” abilities quantitative knowledge, reading and writing ability, fluid reasoning, and “comprehension” knowledge (a subcomponent of which is conceptual understanding). We evaluate the cognitive and developmental alignment of PLMs along these four abilities.
@@ 2.2 Emergent cognitive abilities in Language Models
Recently, the performance of language models has improved as they have increased in size from millions to billions of parameters, trained on larger corpora, and further tuned in novel ways (instruction tuned, RLHF). This has led researchers to increasingly advocate for the use of PLMs as cognitive models Piantadosi (2023); Warstadt and Bowman (2024). Increasing the number of parameters of the models has given rise to emergent abilities that cannot be predicted by extrapolating from the performance of smaller models Wei et al. (2022a). Emergent abilities have been observed in a variety of task types such as multi-task language understanding Hendrycks et al. (2021), grounded conceptual mapping Patel and Pavlick (2022), and truthfulness Lin et al. (2021). In recent work, Hoffmann et al. (2022) and Biderman et al. (2023) have shown the benefits of training a model for more tokens on problem-solving Wei et al. (2022b), common-sense reasoning Sakaguchi et al. (2021), arithmetic abilities Biderman et al. (2023), and linguistic performance Paperno et al. (2016). Although the presence of emergent abilities extends to cognitive science domains Wei et al. (2022b); Goertzel (2023); Hagendorff (2023), prior studies have been piecemeal in their approach and have failed to (1) consider multiple cognitive abilities as specified by theories of psychometric intelligence and (2) move beyond cognitive alignment to also evaluate the developmental alignment of PLMs over training.
2.3 Pre-trained language model use in developmental modeling
Recently, researchers have begun advocating for the use of PLMs for modeling cognitive development in children Kosoy et al. (2023); Salewski et al. (2024). For example, Portelance et al. (2023) and Bhardwaj et al. (2024) suggest the use of language models to predict the age of acquisition of words in children. Researchers have also proposed studying second language acquisition and bilingualism by mapping pre-training steps in PLMs to understand the rate of language development Evanson et al. (2023); Marian (2023); Sharma et al. (2024). We evaluate the proposal that the performance of intermediate training checkpoints of PLMs maps to points during children’s cognitive development.
3 A suite of psychometric intelligence tasks
We assemble a suite of tasks that benchmark PLMs across four facets of psychometric intelligence. Table 1 summarizes the tasks along with the licensing details for public use. The details of each assessment and their respective operationalization are given below.  [^1] We add the tasks to a publically available unified language model testing framework, titled lm-evaluation-harness Gao et al. (2023), to support the evaluation of future models on these psychometric intelligence assessments.
@@ 3.1 Numeric abilities
The question of how humans understand symbolic numbers has been investigated by cognitive scientists for more than half a century. These studies show that people map number symbols to a mental number line (MNL, Figure 2) with a log-compressed psychophysical scale Moyer and Landauer (1967).
Prior research on the numerical abilities of PLMs has focused on higher-order and application-driven aspects of mathematics such arithmetic equations and word problems Burns et al. (2021); Amini et al. (2019); Yuan et al. (2023), exact facts  Lin et al. (2020), and measurement estimation  Zhang et al. (2020). However, these tasks fail to directly tap the foundational cognitive mechanisms underlying human numerical understanding: recruitment of the MNL.
In a recent study, Shah et al. (2023) found evidence for a human-like MNL in various PLMs. They showed that despite lacking explicit neural circuitry to represent numbers, through experience (i.e., vast amounts of training data), PLMs show human-like performance profiles and learn latent representations consistent with the MNL.
We follow Shah et al. (2023) and look for the two behavioral signatures of a compressed MNL representation, the distance effect and the ratio effect. In humans, these are defined as:
  • Distance effect (refer to Figure 1A - top): The greater the distance $|x-y|$ between two numbers $x$ and $y$, the faster they are compared, i.e., the greater (or lesser) number is identified Moyer and Landauer (1967).
  • Ratio effect (refer to Figure 1A - bottom): The time to compare two numbers $x$ and $y$ decreases nonlinearly as a function of the ratio of the larger number over the smaller number $\frac{max(x,y)}{min(x,y)}$ Halberda et al. (2008).
These effects can be mapped to language models by adopting the following linking hypothesis: the greater the cosine similarity of two number representations in a PLM, the more difficult it is to discriminate them (i.e., to judge which one is greater (or lesser)), and thus the longer it takes. While we focus on the Distance and Ratio effects in PLMs, the results for all the effects investigated by Shah et al. (2023) are in Appendix B.1.
Operationalization: We used the same protocol as Shah et al. (2023). For each effect, we test three formats of number representations of PLMs: mixed-case number words, lower-case number words, and digits. We present the $R^{2}$ values for the Distance and Ratio effects, which are averaged across each input representation. The $R^{2}$ values for the distance effect in PLMs are obtained by fitting a linear function predicting the cosine similarity of $x$ and $y$ from their distance $|x-y|$. $R^{2}$ values for the ratio effect in PLMs are obtained by fitting a negative exponential function predicting the normalized cosine similarity of $x$ and $y$ from their ratio $\frac{\max(x,y)}{\min(x,y)}$. Note: This task requires access to the latent representations of models.
@@ 3.2 Linguistic abilities
Language (or verbal) ability is a central component of human cognition and cognitive neuroscience Hagoort (2019). At the dawn of the cognitive revolution, it was conceptualized as a largely innate ability, and language acquisition was thought to require relatively little learning from experience Fodor (1985); Chomsky (2014). More recently, cognitive developmentalists have shown that infants can learn language through exposure to the statistical regularities of the linguistic environment Saffran et al. (1996); Siegelman (2020). These findings have been modeled using multi-layer perceptrons Elman (1996) and, more recently, PLMs Lake and Murphy (2023).
We use BLiMP (Benchmark of Linguistic Minimal Pairs for English) Warstadt et al. (2020) to evaluate the linguistic abilities of each PLM under consideration. BLiMP consists of 67 datasets of 1000 pairs of minimally different sentences which vary in acceptability and span 12 phenomena at three levels of language: morphology, syntax, and semantics. The 12 phenomena are described in Appendix B.2. Each pair consists of one acceptable sentence and one unacceptable sentence which otherwise differ minimally. BLiMP evaluates the models by measuring if they assign a higher probability to the acceptable vs. unacceptable sentence of each pair. Figure 1B shows two examples of minimal pairs.
Operationalization: We use the LM-eval-harness Gao et al. (2023) benchmarking suite to test our models on the BLiMP tasks. We evaluate if a model assigns a higher sequential probability to the acceptable sentence. Note: This requires models that can generate probabilities of tokens.
@@ 3.3 Concept understanding
On encountering a new stimulus, humans categorize it – assign it to a known concept – in order to make inferences about its unobservable properties Murphy (2002). A striking finding is that not all members of a category are equal Rosch (1975). Rather, for a given category (e.g., Bird), some members (e.g., pigeon) are more typical than others (e.g., ostrich). This phenomenon, known as the typicality effect, is a central feature of human categorization Lakoff (2008).
Typicality gradients in human categories can be measured using the production task, where participants are given a category label (e.g., Bird) and asked to list as many members of the category as they can in a limited time Battig and Montague (1969); Van Overschelde et al. (2004); Castro et al. (2021). The typicality of an item is defined as the proportion of participants who produce it.
Language models have shown some evidence of human-like typicality gradients. Heyman and Heyman (2019) used word2vec embeddings to predict the category typicality norms released by De Deyne et al. (2008). More recent work by Misra et al. (2021) and Bhatia and Richie (2022) has looked at correlations of PLMs like BERT, RoBERTa, and GPT-2 to the Rosch (1975) typicality norms for ten categories. Vemuri et al. (2024) performed the most comprehensive study of the alignment of concept understanding in the latent representations of PLMs. We expand upon their task setup to evaluate human-like concept understanding in the PLMs that are the focus here.
@@
  • Surprisal values: For each member of a category, the probability of the sequence a "member" (eg. pigeon) is a "category" (eg. bird). This method requires access to the probability of each token in a sequence.
  • Prompting: Prompt the models with the following design: Guidelines, Query, and Options. The Guideline highlights the task of re-ranking the members given in the Options based on appropriateness with the Query. The Query consists of the in-filling task: A   is a [category name]. The Options are each of the possible members of the category. Given the complexity of the prompting, usable outputs are only obtained from models that are larger than 30 billion parameters.